OpenStudy (anonymous):

Solve for X, be sure to show all WORK ! x-8 1 --------- = ------ X+4 5

OpenStudy (anonymous):

are the X's supposed to be different sizes??

OpenStudy (anonymous):

No . That was a accident that i made them diffrent sizes .

OpenStudy (anonymous):

ok... Another way to write it: \[x+4*1 = x-8*5\]

OpenStudy (anonymous):

\[\frac{ x - 8 }{ x - 4 } = \frac{ 1 }{ 5 }\] This is a proportion, so we can multiply the diagonals and set them equal to each other. \[5(x-8) = x-4\] Distribute: \[5x - 40= x-4\] Simplify it a little: \[4x - 40= -4\] Now I'll leave you to solve the rest.

OpenStudy (anonymous):

1(x+4) = 5(x-8) 1x+4 = 5x-40 -1x -1x 4 = 4x-40 +40 +40 44 = 4x 44/4 = 4x/4 44/4 = x x = 11

OpenStudy (anonymous):

@Forever_Me

OpenStudy (anonymous):

Do u see wat i did?

OpenStudy (anonymous):

Yes i see . THANK YOU SOOOOOOOOOOOO MUCH !!! I appreciate it :) @LizzyLove<3

OpenStudy (anonymous):

Your very Welcome ;)

OpenStudy (anonymous):

Do you need anymore help?? @Forever_Me

OpenStudy (anonymous):

Yes @LizzyLove<3 x-4 --------- <4 x+5

OpenStudy (anonymous):

Change '<' to an equal sign... then write four as 4/1

OpenStudy (anonymous):

1(x-4) = 4(x+5) 1x-4 = 4x + 20 -1x -1x -4 = 3x + 20 -20 -20 -24 = 3x -24/3 = 3x/3 -8 = x

OpenStudy (anonymous):

@Forever_Me